OpenStudy (anonymous):

how do I solve this problem. The difference of two numbers is 67. Larger number is 3 less than 6 times the smaller. Find the two numbers.

OpenStudy (johnweldon1993):

Set up a system of equations... you know that x - y = 67 and lets say larger number is y...so y = 6x - 3 since the larger number is 3 less...than 6 times the smaller so your 2 equations are x - y = 67 y = 6x - 3 since you now know what y = just plug that in for y in the first equation x - (6x - 3) = 67 can you solve for x?

OpenStudy (anonymous):

Thank you very Much :) I was not sure how to set it up!

OpenStudy (johnweldon1993):

Anytime! :)